cookiemonster:

Part A: The area of a square is (9x^2 24x 16) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. (5 points) Part B: The area of a rectangle is (16x^2 − 25y^2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. (5 points)

Elsa213:

Brainliest? >.>

cookiemonster:

hehe

cookiemonster:

@Elsa213 my mistake sorry

cookiemonster:

@Elsa213 can you help me with the question please

Elsa213:

``` 9x^2 + 24x + 16 = (3x + 4)(3x + 4) length of each side = 3x + 4 units. Par B 16x^2 - 25y^2 = (4x + 5y )( 4x - 5y) length = 4x + 5y and width = 4x - 5y ``` Source: Google

cookiemonster:

thank you so much! :)

cookiemonster:

i have one more question :)

cookiemonster:

An expression is shown below: f(x) = 5x^2 + 2x − 3 Part A: What are the x-intercepts of the graph of f(x)? Show your work. (2 points) Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points) Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points)

cookiemonster:

@Elsa213

Mehek:

Part A: To find x intercepts, first factor 5x^2 + 2x − 3 (5x - 3)(x + 1) Solve for x 5x - 3 = 0 5x = 3 x = 3/5 x + 1 = 0 x = -1 So 3/5 and -1

cookiemonster:

ok awesome thats part a :)

Mehek:

Yes now since the first term is positive, that means the graph is going in a U shape and it has a minimum

Mehek:

Which is at (-0.2, -3.2)

Mehek:

To show the work to find x it is x = -b/2a so x = -2/2*5 x = -0.2 Plug that in for x and you get y = -3.2

Mehek:

Part C, to graph you would plot the x intercepts and minimum and then if you want more points just put any number for x, find the y and plot

cookiemonster:

ok so wait for part b, its..?

cookiemonster:

@Mehek

Mehek:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek Yes now since the first term is positive, that means the graph is going in a U shape and it has a minimum \(\color{#0cbb34}{\text{End of Quote}}\)

Mehek:

And then everything after that before the "part C"

cookiemonster:

ok perfect thank you so much